# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Agent Systems

### v1

@@ Abstract
Precise localization of multi-agent systems (MAS) in harsh environments is a critical challenge for swarm applications, and cooperative localization is considered a key solution to this issue. Among all solutions, distributed cooperative localization (DCL) has garnered widespread attention due to its robustness and scalability. The main challenge of DCL lies in how to fuse relative measurements between agents under unknown correlations. To address this, covariance intersection (CI) was introduced to DCL. However, the classical CI optimization criteria suffer from issues such as scale imbalance and correlation mismatch during the fusion process. These deficiencies are not as pronounced in 2D scenarios, where the state space is relatively simple and the observability of each state component is well. However, in 3D scenarios, where the state space is more complex and there are significant disparities in the scale and observability of state components, performance degradation becomes severe. This necessitates the design of specialized mechanisms to improve the data fusion process. In this paper, we identify three main drawbacks of the classical CI optimization criteria in recursive filtering and introduce a weighting mechanism, namely weighted covariance intersection (WCI), to improve its performance. We then introduce WCI into range-based distributed cooperative localization in 3D scenarios, developing a concurrent fusion strategy for multiple distance measurements and designing a weighting matrix based on the error propagation rule of the inertial navigation system (INS). Simulation results demonstrate that the proposed WCI significantly enhances cooperative localization performance compared to classical CI, while the distributed approach outperforms the centralized approach in terms of robustness, scalability, and is more suitable for large-scale swarms.
Index Terms:
Distributed cooperative localization, multi-agent systems, correlation, covariance intersection, weighting mechanism.
@@ I Introduction
Intelligent agents, represented by unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs), have found wide applications in various fields such as environmental monitoring, disaster rescue, and military operations [1, 2, 3]. In recent years, multi-agent systems (MAS) have gained significant attention due to their high flexibility, collaboration capabilities, and robustness, enabling them to efficiently perform complex tasks [4, 5]. The evolution from individual agent intelligence to swarm intelligence has become a key trend in the development of unmanned technologies. Accurate positioning services are the foundation for MAS to execute tasks. Currently, the Global Navigation Satellite System (GNSS) is one of the most widely used positioning methods in MAS [6]. Although GNSS provides all-day and all-weather positioning services in outdoor environments, its signals can be severely obstructed and become unavailable in indoor and underwater scenarios [7]. Moreover, GNSS is not always reliable and is susceptible to interference and spoofing [8]. In contrast, cooperative localization, as an alternative solution, can effectively improve positioning accuracy and availability through relative measurements, information exchange, and data fusion between multiple agents [9].
@@
Since the concept of cooperative localization was first introduced by [10], it has attracted widespread attention and made significant progress over the past three decades. Cooperative localization relies on relative measurements between agents, but these measurements may be correlated with the agents’ states, which must be accounted for during state updates. We refer to the state update process using correlated measurements as correlated update in later illustration. Depending on whether the computation occurs at a central unit or locally at each agent, cooperative localization can be divided into centralized cooperative localization (CCL) and distributed cooperative localization (DCL) [11]. Typically, the two schemes tackle the constantly changing state correlations between agents in different ways. CCL tracks the state correlations between agents with the requirement that all data be collected at a central processing unit, leading to significant communication overhead and poor scalability and robustness. In contrast, DCL typically neither requires nor is capable of tracking state correlations between agents, making it naturally more suitable for large-scale swarms and complex scenarios. However, this also presents the primary challenge of how to perform effective correlated update under unknown state correlations between agents.
@@
Covariance intersection (CI) was initially proposed as a distributed data fusion method capable of producing a consistent and conservative estimate under unknown correlations between data sources [12], and has since been extensively applied to recursive filtering tasks involving unknown correlations between measurements and estimated states, such as target tracking and cooperative perception [13, 14]. Since Arambel first introduced CI to cooperative localization [15], CI and its derivatives have been widely applied in DCL with great success. Carrillo and Li used relative pose observations to obtain two estimates of the agent’s pose and fused them using CI and split covariance intersection (SCI), respectively [11, 16, 17]. Pierre extended CI-based correlated updates to partial observation scenarios, achieving DCL based solely on ranging measurements [18]. Chang also utilized CI to develop a cooperative localization scheme that decouples measurement and communication, improving the robustness and scalability of distributed cooperative localization [19]. However, the above CI-based cooperative localization algorithms are all designed for two-dimensional (2D) scenarios. There is still limited research on DCL for MAS, such as UAV swarms, moving in three-dimensional (3D) environments.
The performance of CI depends on the selection of the weight parameter $\omega$, which is influenced by the optimization criterion. Previous studies have primarily treated CI as a plug-in method for cooperative localization, directly applying classical CI optimization criteria, such as trace or determinant-based metrics of the fused covariance matrix, without considering how different optimization criteria affect the performance of cooperative localization tasks. Furthermore, unlike Kalman filtering (KF), which is inherently a recursive filtering method, CI is a distributed data fusion technique focused on providing consistent estimates in a single-step fusion, and thus exhibits inherent limitations when applied within recursive filtering [20].We investigate the impact of different optimization criteria on cooperative localization performance and identify three major drawbacks of classical CI: scale imbalance, correlation mismatch, and lack of task relevance. Among these, the first two drawbacks are particularly critical, as they can significantly degrade state estimation performance especially when state vector components have large disparities in scale and observability. A detailed discussion of these drawbacks is provided in Section IV. These deficiencies are less pronounced in 2D scenarios, since the state vector consists only of 2D position and orientation, and the scale disparity is relatively small, and the observability of each state component is well. However, in 3D scenarios, where the state space is more complex and there are significant disparities in the scale and observability of state components, performance degradation becomes pronounced. In particular, trace-based optimization in CI performs poorly for small-scale state components, such as attitude and IMU biases, while determinant-based optimization of CI leads to degraded accuracy in position estimation. This highlights the need for specialized mechanisms to improve the correlated update process for DCL in 3D scenarios.
@@
The weighting mechanism is a widely employed method for controlling the data fusion process [21]. Its primary advantage lies in its flexibility and controllability, enabling the adjustment of weights based on specific requirements to achieve optimal fusion results. By incorporating this weighting mechanism into CI, we can improve its performance in correlated data fusion, particularly in balancing state components with significant scale disparities. Furthermore, the weighting mechanism provides a promising strategy for making the optimization criterion task-oriented. In practical applications, different state components can have varying impacts on the final result, and general MSE optimization often fails to accurately reflect the actual contribution of the estimated states to the task at hand. In contrast, weighted mean squared error (WMSE) optimization offers clearer and more relevant significance in these scenarios [22]. Although the concept of weighted covariance intersection (WCI) was first introduced in [15], it primarily focused on enhancing the estimation performance of specific state subspaces and did not highlight the necessity of WCI in scenarios with significant disparities in the scale of state components. Furthermore, it lacked an interpretation of the optimization significance associated with WCI. These two limitations have, to some extent, restricted its application in recursive filtering tasks, such as cooperative localization. In contrast, in this paper, we carefully address these two challenges and successfully integrate WCI into DCL for 3D scenarios, achieving superior performance compared to classical CI.
The main contributions of this paper are summarized as follows:
  • We identify three major drawbacks of CI in recursive filtering: correlation mismatch, scale imbalance, and lack of task relevance. These drawbacks present potential risks in recursive filtering that must be carefully considered and addressed, especially when there are significant differences in the scale and observability of state components. We analyze how WCI can potentially mitigate these drawbacks and clarify its physical meaning in the context of WMSE, as well as its properties.
  • We introduce WCI into range-based distributed cooperative localization in 3D scenarios. We develop a concurrent fusion strategy for multiple distance measurements and design a weighting matrix according to the error propagation rule of INS, addressing the issue of unbalanced and degraded estimation performance between different states in classical CI for this task.
  • We conduct extensive simulation experiments to validate the effectiveness and superiority of the proposed methods. The results show that although WCI does not completely eliminate the inherent drawbacks of CI, it significantly reduces their impact, leading to improved state estimation accuracy in cooperative localization tasks.
@@
The remainder of this paper is organized as follows. In Section II, we introduce the setup of the MAS under study, including the system and state composition, as well as range measurement models. In Section III, we introduce the overall distributed cooperative localization framework and focus our work on improving the correlated update process. Section IV elaborates on the proposed WCI-based correlated update process for DCL in 3D scenarios, including a well-designed weighting matrix selection strategy based on the error propagation rule of INS. In Section V, we present extensive simulations to validate the effectiveness of the proposed DCL method and demonstrate its superiority over CCL method. Finally, the conclusions are summarized in Section VII.
@@ II System Setting
We consider a range-based cooperative localization system consisting of a MAS with $N_{u}$ agents moving in 3D space and $N_{a}$ fixed anchors with known positions, as shown in Fig. 1. Since each agent may not be able to connect to a sufficient number of anchors due to communication range limitations or obstructions, it relies on cooperation between agents to achieve accurate state estimation. Each agent is equipped with an onboard IMU to capture its self-motion. In addition, both agents and anchors are equipped with rigidly attached ranging modules. Through signal transceiving, each agent can measure the relative distances to other agents and anchors within its communication range [^1] In this study, anchors are assumed to function solely as signal transmitters, with measurements occurring exclusively on the agent side.. Furthermore, agents and anchors are equipped with communication modules for exchanging information.
Since we use IMU as the proprioceptive sensor, the state of agent $i$ ($i=1,\ldots,N_{a}$) is defined as follows:
Figure 1: Schematic diagram of a range-based cooperative localization system. Due to communication range limitations or obstructions, agents may not be able to obtain sufficient distance measurements with anchors for state estimation. However, through cooperation between agents (including relative distance measurements, information exchange and data fusion), agents can achieve accurate state estimation in such cases.
@@
where $\prescript{I}{}{\boldsymbol{p}}_{i}^{n}$, $\prescript{I}{}{\boldsymbol{v}}_{i}^{n}$ and $\boldsymbol{q}_{i}^{n}$ represents the position, velocity and orientation (expressed in quaternion form) of the agent in the navigation frame $\mathcal{N}$, respectively. $\boldsymbol{b\scriptstyle g}_{i}$ and $\boldsymbol{b\scriptstyle a}_{i}$ represent the biases of the gyroscope and the accelerometer of the IMU, respectively. The superscript $I$ on the left indicates that the state corresponds to the position of the IMU on the agent. In IMU-based state estimation, system error propagation and state updates are typically performed using the error-state formulation, rather than directly using the nominal state [23]. This approach mitigates the redundancy inherent in quaternion-based attitude representation and addresses the challenges associated with nonlinearity. We define the error-state of agent $i$ as follows:
@@
In general, the IMU and the ranging module are not co-located on the agent, and there is a lever arm between them. Assuming that each agent has an identical lever arm from the IMU to the ranging module, denoted as $\boldsymbol{l}^{b}$ in the agent’s body-fixed frame $\mathcal{B}$. The relationship between the IMU’s position and the ranging module’s position of agent $i$ can be expressed as follows:
@@
$d_{ij}^{\mathcal{U}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\rVert\qquad d_{im}^{\mathcal{A}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}_{m}^{n}}\rVert$ (4)
where $d_{ij}^{\mathcal{U}}$ is the distance between the ranging modules of agent $i$ and agent $j$, while $d_{im}^{\mathcal{A}}$ is the distance between the ranging modules of agent $i$ and anchor $m$ ($\prescript{A}{}{\boldsymbol{p}_{m}^{n}}$ is the position of anchor $m$ expressed in frame $\mathcal{N}$.). The superscripts $\mathcal{U}$ and $\mathcal{A}$ are used to distinguish the relative distances between two agents and those between an agent and an anchor. The distance measurements can be modeled as
${\tilde{d}_{ij}^{\mathcal{U}}}={{d}_{ij}^{\mathcal{U}}}+n_{d_{ij}^{\mathcal{U}}}\qquad{\tilde{d}_{im}^{\mathcal{A}}}={{d}_{im}^{\mathcal{A}}}+n_{d_{im}^{\mathcal{A}}}$ (5)
where $n_{d_{ij}^{\mathcal{U}}}$ and $n_{d_{im}^{\mathcal{A}}}$ are assumed to be i.i.d zero-mean Gaussian with variance $\sigma_{r}^{2}$.
@@ III Distributed Cooperative Localization Framework
In this section, we introduce a range-based distributed cooperative localization framework for 3D scenarios. As illustrated in Fig. 2, we use the state estimation process of agent $i$ at time step $k$ as an example to explain the procedure of the general DCL framework.
In distributed cooperative localization, each agent only estimates its own state and maintains its own state covariance matrix. The state estimation of an agent can be divided into two phases: state prediction and state update. In the state prediction phase, the agent uses its onboard IMU outputs to predict state changes and adjust the covariance matrix for error propagation. In the state update phase, the agent uses distance measurements to other agents and anchors, along with information exchanged with other agents, to correct its state through data fusion methods. More specifically, distance measurements can be categorized into two types: measurements between agents and anchors, and measurements between agents. The former’s noise is independent of agents’ state, while the latter’s is correlated with agents’ state. Therefore, different data fusion methods are required for state updates.

### v2

# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Vehicle Systems

@@ Abstract
Cooperative localization is considered a key solution for enabling autonomous navigation of multi-vehicle systems (MVS) in GNSS-denied environments. Among all solutions, distributed cooperative localization (DCL) has garnered widespread attention due to its robustness and scalability, making it well-suited for large-scale MVS. To address the challenge of untrackable state correlations between vehicles in a distributed framework, covariance intersection (CI) has been introduced as a means to fuse relative measurements under unknown correlations. However, existing studies treat CI merely as a plug-in method, applying traditional optimization criteria directly and focusing only on simple two-dimensional (2D) scenarios. When directly extended to three-dimensional (3D) scenarios with more complex state space (higher dimensions, additional state components, and significant disparities in scale and observability among state components), traditional methods fail to achieve balanced state estimation across all state components, leading to a significant degradation in the estimation accuracy of some state components. This highlights the need to design specialized mechanisms to improve the data fusion process. In this paper, we introduce a weighting mechanism, namely the weighted covariance intersection (WCI), to regulate the fusion process of CI. A concurrent fusion strategy for multiple distance measurements and a dedicated weighting matrix based on the error propagation rule of the inertial navigation system (INS) are developed for the data fusion process in DCL. Simulation results demonstrate that the proposed WCI significantly enhances cooperative localization performance compared to traditional CI, while the distributed approach outperforms the centralized approach in terms of robustness and scalability.
@@ I Introduction
Autonomous vehicles, represented by unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs), have found wide applications in various fields such as environmental monitoring, disaster rescue, and military operations [1, 2, 3]. In recent years, multi-vehicle systems (MVS) have gained significant attention due to their high flexibility, collaboration capabilities, and robustness, allowing them to efficiently perform complex tasks [4, 5]. Accurate positioning services are the foundation for MVS to perform tasks. Currently, the Global Navigation Satellite System (GNSS) is one of the most widely used positioning methods in MVS [6]. Although GNSS provides all-day and all-weather positioning services in outdoor environments, its signals can be severely obstructed and become unavailable in indoor and underwater scenarios[7]. Local positioning systems (LPSs) require the prior deployment of infrastructure[8]. In contrast, cooperative localization, as an alternative solution, can effectively improve positioning accuracy and availability in GNSS-denied environments through vehicle collaboration[9, 10, 11].
@@
Since the concept of cooperative localization was first introduced by [12], it has attracted widespread attention and made significant progress over the past three decades. Typically, cooperative localization involves relative measurements between vehicles, vehicle-to-vehicle communication (V2V) for information exchange, and the fusion of data of local and cooperative estimates[13], where well-localized vehicles can assist others to achieve better global localization performance [14]. Among various solutions, filter-based cooperative localization, which tracks vehicle poses (including both position and attitude) through recursive filtering, is particularly effective in enhancing autonomous navigation capabilities. Since each vehicle utilizes the state information of other vehicles to estimate its own state, their states become correlated, which must be accounted for during the filtering process when updating states. We refer to the state update process using measurements between vehicles as correlated update in later illustration. Depending on whether the computation occurs at a central unit or locally at each vehicle, cooperative localization can be divided into centralized cooperative localization (CCL) and distributed cooperative localization (DCL) [15]. Typically, the two schemes tackle the constantly changing state correlations between vehicles in different ways. CCL tracks the state correlations between vehicles with the requirement that all data be collected at a central processing unit, leading to significant communication overhead and poor scalability and robustness. In contrast, DCL typically neither requires nor is capable of tracking state correlations between vehicles, making it naturally more suitable for large-scale swarms and complex scenarios. However, this also presents the primary challenge of how to perform effective correlated update under unknown state correlations between vehicles.
@@
Covariance intersection (CI) was initially proposed as a distributed data fusion method capable of producing a consistent and conservative estimate under unknown correlations between data sources [16], and has been extensively applied to recursive filtering tasks involving unknown correlations between measurements and estimated states, such as target tracking and cooperative perception [17, 18]. Since Arambel first introduced CI to cooperative localization [19], CI and its derivatives have been widely applied in cooperative vehicle localization with great success. Carrillo and Li used relative pose observations to obtain two estimates of the vehicle’s pose and fused them using CI and split covariance intersection (SCI), respectively [15, 20, 21]. Pierre extended CI-based correlated updates to partial observation scenarios, achieving DCL based solely on distance measurements [22]. Chang also utilized CI to develop a cooperative localization scheme that decouples measurement and communication, improving the robustness and scalability of distributed cooperative localization [23]. Fang considered the problem of cooperative localization under data sources of various error types and used the adaptive cubature split covariance intersection filter to estimate the vehicles’ states [24].
Despite the extensive research on CI-based cooperative localization, two major limitations remain. First, existing studies treat CI merely as a plug-in data fusion technique, directly applying traditional trace-based [21, 23] or determinant-based [15, 20, 22, 24] metrics of the fused covariance matrix, without considering how different optimization criteria affect the performance of cooperative localization tasks. Second, these studies are all designed for MVS moving in two-dimensional (2D) scenarios, where vehicles adopt unicycle model with a state space limited to three dimensions. There is still limited research on DCL for MVS, such as UAV swarms, operating in three-dimensional (3D) environments. Compared with 2D scenarios, the state space of vehicles is more complex, with higher dimensions, additional state components, and significant disparities in the scale and observability of state components. If existing approaches are directly applied to 3D vehicle cooperative localization, due to significant disparities in the scale and observability of state components, the CI fusion process under traditional CI criteria is dominated by position, and the uncertainty of state components with small scales and poor observability such as attitude will inflate greatly and their estimation performance degrades severely, as shown in Fig. 1. It is important to note that not only is attitude crucial in the operation of vehicles, but the estimation accuracy of small-scale states, such as attitude, also significantly impacts the long-term stability of localization. If the estimation errors of such states are large, the autonomous navigation capabilities of the vehicles will be severely affected, and the position error become highly susceptible to divergence, especially in situations without sufficient observation. We expect an improved CI optimization criterion that can balance and improve the estimation of all state components. As a result, specialized mechanisms must be designed to regulate and improve the data fusion process of CI in 3D cooperative localization task.
@@
The weighting mechanism is widely employed to control the data fusion process [25]. Its primary advantage lies in its flexibility and controllability, enabling the adjustment of weights based on specific requirements to achieve optimal fusion results. By incorporating this weighting mechanism into CI, it’s promising to improve its performance in correlated data fusion, particularly in balancing state components with significant scale and observability disparities. Furthermore, the weighting mechanism provides a strategy to make the optimization criterion task-oriented. In practical applications, different state components can have varying impacts on the final result, and general mean squared error (MSE) optimization often fails to accurately reflect the actual contribution of the estimated states to the task at hand. In contrast, weighted mean squared error (WMSE) optimization offers clearer and more relevant significance in these scenarios [26]. Although the concept of weighted covariance intersection (WCI) was first introduced in [19], it primarily focused on enhancing the estimation performance of specific state subspaces and did not highlight the necessity of WCI in scenarios with significant disparities in the scale of state components. Furthermore, it lacked an interpretation of the optimization significance associated with WCI. These two limitations have, to some extent, restricted its application in recursive filtering tasks, such as cooperative localization. In contrast, in this paper, we carefully address these two challenges and successfully integrate WCI into DCL for 3D scenarios, achieving superior performance compared to traditional CI.
Figure 1: The influence of traditional CI criteria on 3D cooperative localization. For simplicity of illustration, we select the subspace of its whole state space with 2D position and attitude only.
In this paper, to achieve balanced estimation of different state components, we introduce WCI into range-based distributed cooperative localization in 3D scenarios. To improve the data fusion process, we develop a concurrent fusion strategy for multiple distance measurements that adapts to the characteristics of CI fusion and design a dedicated weighting matrix based on the error propagation rule of the inertial navigation system (INS). Extensive simulation experiments are conducted to validate the effectiveness and superiority of the proposed methods. The results show that WCI can effectively improve and balance state estimation in 3D cooperative localization tasks compared to traditional CI, and the distributed approach outperforms the centralized approach in terms of robustness and scalability.
@@
The remainder of this paper is organized as follows. In Section II, we introduce the setup of the MVS under study, including the system and state composition, as well as range measurement models. In Section III, we introduce the overall distributed cooperative localization framework and focus our work on the correlated update process. Section IV elaborates on the proposed WCI-based correlated update process for DCL in 3D scenarios, including a concurrent fusion strategy for multiple distance measurements and a well-designed weighting matrix based on the error propagation rule of INS. In Section V, we present extensive simulations to validate the effectiveness of the proposed DCL method and demonstrate its superiority over CCL method. Finally, the conclusions are summarized in Section VII.
@@ II System Setting
We consider a range-based cooperative localization system consisting of a MVS with $N_{v}$ vehicles moving in 3D space and $N_{a}$ fixed anchors with known positions, as shown in Fig. 2. Each vehicle is equipped with an onboard inertial measurement unit (IMU) to capture its self-motion. In addition, both vehicles and anchors are equipped with rigidly attached ranging modules. Through signal transceiving, each vehicle can measure the relative distances to other vehicles and anchors within its communication range [^1] In this study, anchors are assumed to function solely as signal transmitters, with measurements occurring exclusively on the vehicle side.. Furthermore, vehicles and anchors are equipped with communication modules for information exchange.
Since we use IMU as the proprioceptive sensor, the state of vehicle $i$ ($i=1,\ldots,N_{v}$) is defined as follows:
Figure 2: Schematic diagram of a range-based cooperative localization system. Due to communication range limitations or obstructions, vehicles may not be able to obtain sufficient distance measurements with anchors for state estimation. However, through vehicle collaboration, all vehicles can achieve accurate state estimation in such cases.
@@
where $\prescript{I}{}{\boldsymbol{p}}_{i}^{n}$, $\prescript{I}{}{\boldsymbol{v}}_{i}^{n}$ and $\boldsymbol{q}_{i}^{n}$ represents the position, velocity and attitude (expressed in quaternion form) of the vehicle in the navigation frame $\mathcal{N}$, respectively. $\boldsymbol{b\scriptstyle g}_{i}$ and $\boldsymbol{b\scriptstyle a}_{i}$ represent the biases of the gyroscope and the accelerometer of the IMU, respectively. The superscript $I$ on the left indicates that the state corresponds to the position of the IMU on the vehicle. In IMU-based state estimation, system error propagation and state updates are typically performed using the error-state formulation rather than directly using the nominal state [27], which mitigates the redundancy inherent in quaternion-based attitude representation and addresses the challenges associated with nonlinearity. We define the error-state of vehicle $i$ as follows:
@@
In general, the IMU and the ranging module are not co-located on the vehicle, and there is a lever arm between them. Assuming that each vehicle has an identical lever arm from the IMU to the ranging module, denoted as $\boldsymbol{l}^{b}$ in the vehicle’s body-fixed frame $\mathcal{B}$. The relationship between the IMU’s position and the ranging module’s position of vehicle $i$ can be expressed as follows:
@@
$d_{ij}^{\mathcal{V}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\rVert\qquad d_{im}^{\mathcal{A}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}_{m}^{n}}\rVert$ (4)
where $d_{ij}^{\mathcal{V}}$ is the distance between the ranging modules of vehicle $i$ and vehicle $j$, while $d_{im}^{\mathcal{A}}$ is the distance between the ranging modules of vehicle $i$ and anchor $m$ ($\prescript{A}{}{\boldsymbol{p}_{m}^{n}}$ is the position of anchor $m$ expressed in frame $\mathcal{N}$.). The superscripts $\mathcal{V}$ and $\mathcal{A}$ are used to distinguish the relative distances between two vehicles and those between an vehicle and an anchor. The distance measurements can be modeled as
${\tilde{d}_{ij}^{\mathcal{V}}}={{d}_{ij}^{\mathcal{V}}}+n_{d_{ij}^{\mathcal{V}}}\qquad{\tilde{d}_{im}^{\mathcal{A}}}={{d}_{im}^{\mathcal{A}}}+n_{d_{im}^{\mathcal{A}}}$ (5)
where $n_{d_{ij}^{\mathcal{V}}}$ and $n_{d_{im}^{\mathcal{A}}}$ are assumed to be i.i.d zero-mean Gaussian with variance $\sigma_{r}^{2}$.
@@ III Distributed Cooperative Localization Framework
In this section, we introduce a range-based distributed cooperative localization framework for 3D scenarios. As illustrated in Fig. 3, we use the state estimation process of vehicle $i$ at time step $k$ as an example to explain the procedure of the general DCL framework.
In distributed cooperative localization, each vehicle only estimates its own state and maintains its own state covariance matrix. The state estimation of an vehicle can be divided into two phases: state prediction and state update. In the state prediction phase, the vehicle uses its onboard IMU outputs to predict state changes and adjust the covariance matrix for error propagation. In the state update phase, the vehicle uses distance measurements to other vehicles and anchors, along with information exchanged with other vehicles, to correct its state through data fusion. More specifically, distance measurements can be categorized into two types based on the object being measured: measurements between vehicles and anchors, and measurements between vehicles. Different data fusion methods are required for the two types of measurements.